Circular No. 6697

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SATURN C. Roddier, F. Roddier, J. E. Graves and M. J. Northcott, Institute for

Astronomy, University of Hawaii (UH), report: “Infrared images of Saturn’s rings were taken on 1995 Aug. 12 with the UH adaptive-optics system mounted on the Canada-France-Hawaii telescope. After careful processing, evidence was found for a faint streak of light moving away from Saturn at the level of the E ring. The object is elongated in the direction of Saturn with a sharp edge on Saturn’s side and a fuzzy tail on the other side. Assuming motion on a keplerian orbit, the object’s orbital radius must be within a few km of that of Saturn II (Enceladus). The object’s sharp edge is ahead of Enceladus by an orbital longitude of 75^{o}. Its orbit is inclined by ∼ 2^{o} with respect to the ring plane. The object could possibly be an arc structure confined by gravitational interaction with both Saturn I (Mimas) and Enceladus.”

GRB 970616 J. C. Wheeler, University of Texas, reports that R-band images cen-

tered on the location of GRB 970616 (IAUC 6683) were made by F. Ma, Z. Shang, A. Howell, P. Rathbun and Y. Sheffer with the McDonald Observatory’s 0.76-m telescope (field of view 46′.2 × 46′.2) on June 19.4, 20.4, 21.4, 22.4, 23.4 and 25.4 UT. No variables or galaxies were seen down to R ∼ 20 ± 0.5 (limited by moonlight) in the error box of the x-ray variable reported by Murakami et al. (IAUC 6687). The first of these optical scans was made only 64 hours after the GRB occurred.

NOVA SCORPII 1997 Visual magnitude estimates: June 10.20 UT, 10.4 (W. G. Dillon, Mis-

souri City, TX); 12.10, 11.7 (A. Pereira, Cabo da Roca, Portugal); 14.99, 11.3 (D. Rodriguez, Madrid, Spain); 28.18, 12.4 (Dillon).

COMET C/1997 N1 (TABUR) Total visual magnitude estimates: July 6.92 UT, 10.0 (A. Pearce, Subi-

aco, Western Australia, 0.2-m reflector); 7.80, 10.3 (D. A. J. Seargent, The Entrance, N.S.W., 25 × 100 binoculars); 8.91, 10.4 (Pearce).

1997 July 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden